Leon Jackson invests $48,800 at 8% annual interest, leaving the money invested without withdrawing any of the interest for 8 years. At the end of the 8 years, Leon withdraws the accumulated amount of money.

New attempt is in progress. Some of the new entries may impact the last attempt grading.Your answer is incorrect.

Compute the amount Leon would withdraw assuming the investment earns simple interest. (Round answers to 0 decimal places, e.g. 458,581.)

Total withdrawn

$enter total withdrawn in dollars rounded to 0 decimal places

eTextbook and Media

New attempt is in progress. Some of the new entries may impact the last attempt grading.Your answer is incorrect.

Click here to view factor tables

Compute the amount Leon would withdraw assuming the investment earns interest compounded annually. (Round factor values to 5 decimal places, e.g. 1.25124 and final answer to 0 decimal places, e.g. 458,581.)

Total withdrawn

$enter total withdrawn in dollars rounded to 0 decimal places

New attempt is in progress. Some of the new entries may impact the last attempt grading.Your answer is incorrect.

Click here to view factor tables

Compute the amount Leon would withdraw assuming the investment earns interest compounded semiannually. (Round factor values to 5 decimal places, e.g. 1.25124 and final answer to 0 decimal places, e.g. 458,581.)

Total withdrawn

$enter total withdrawn in dollars rounded to 0 decimal places

eText

eTextbook and Mediabook and Media
Answer #1

Solution 1:

Total withdrawn amount = Principal + Interest = $48,800 + $48,800*8%8* = $80,032

Solution 2:

Interest rate = 8 %

Nos of periods = 8 years

Amount withdrawn = $48,800 * (1+0.08)^8 = $48,800*1.85093 = $90,325

Solution 3:

Interest rate = 8 %, 4% semiannual

Nos of periods = 8 years, 16 semiannual period

Amount withdrawn = $48,800 * (1+0.04)^16 = $48,800*1.87298 = $91,401